A compound was found to contain 49.98% C, and 10.47% H. The molar mass of the compound is 58.12 g/mol. Determine the molecular formula.


1
Expert's answer
2022-01-24T13:30:02-0500

Data Provided: Composition of the compound

Carbon – 49.98%

Hydrogen- 10.47%

Molecular mass- 58.12 g/mol

Required –Molecular formula

Step 1: We change each percentage to an expression of the mass of each element

That is 49.98% C becomes 49.98g C and 10.47 % H becomes 10.47g H

Step 2: We convert amount of each element in grams to its moles.

Moles of C = 49.98g C × 1 mole C/12.01g C= 4.1615 moles

Moles of H = 10.47g H × 1 mole H/1.008g H = 10.3869 moles

Each mole value is then divided by the smallest mole value:

4.1615 moles/4.1615 moles= 1×2= 2

10.3869 moles /4.1615 moles = 2.4959×2=4.99195 =5 (We multiply by 2 to convert into in integer)

Thus the empirical formula of the compound is C2H5.

Using the relationship:

 (Empirical formula mass g/mol) n = molecular mass g/mol

Where n is a whole number multiple.

 n = 58.12 g/mole/(2×12.01 + 5×1.008) g/mol

n = 2

 Molecular formula = (Empirical formula) n

Thus, the molecular formula is (C2H5)2 = C4H10